# Pretraining with random noise for uncertainty calibration

### v1

@@ Results
We investigated whether the pre-calibration effect observed with random noise training extends to unseen, more realistic data. Specifically, we examined how the network’s confidence on CIFAR-10 and SVHN 54 datasets changes during random noise pretraining (Fig. 4i, Left) by measuring the network’s predictions for real data after each epoch during the pretraining phase. Initially, we observed that the confidence for unseen CIFAR-10 data was well-calibrated and approached the chance level, while accuracy remained at the chance level throughout random noise training (Fig. 4i, Right). We confirmed that this trend was consistent across different unseen datasets, including CIFAR-10 (Fig. 4j, CIFAR-10; w/ vs. w/o random pretraining, $n_{\text{trial}}=10000$, Wilcoxon rank-sum test, $P<10^{-3}$) and SVHN (Fig. 4j, SVHN; w/ vs. w/o random pretraining, $n_{\text{trial}}=10000$, Wilcoxon rank-sum test, $P<10^{-3}$). While accuracy remained at the chance level, the network maximized uncertainty for unseen data, aligning confidence with accuracy at the chance level. As a result, calibration error — the discrepancy between confidence and accuracy — was significantly reduced for unseen data (Fig. 4j, w/ vs. w/o random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, CIFAR-10, $P<10^{-3}$; SVHN, $P<10^{-3}$). These findings suggest that networks initialized conventionally tend to exhibit overconfidence and class bias even before training on data. In contrast, random noise pretraining effectively eliminates this initial bias and calibrates confidence to the chance level, maximizing uncertainty for unseen datasets. This result indicates that random noise pretraining can serve as an novel initialization strategy, offering an alternative to conventional random initialization 53.
Aligning confidence levels with actual accuracy during random noise pretraining
Fig. 5: Pre-calibration enables learning with matching confidence and accuracy
@@
(a) Visualization of loss-accuracy duality. Each dot represents the test loss and accuracy at each epoch of training. Two points are selected for comparison of calibration: selected networks achieving the same accuracy for both the random noise pretrained network and the data-only training network (1, 2), as well as fully converged networks trained for the same number of epochs (3, 4). (b-c) Comparison between random noise pretrained networks and data-only training networks. (Left) Reliability diagram, showing the relationship between predicted confidence and accuracy. (Right) Calibration error, measured as Expected Calibration Error (ECE). (b) Accuracy-controlled comparison (1 vs. 2). (c) Epoch-controlled comparison (3 vs. 4). (d) Scatter plot of confidence versus accuracy for the training dataset in a data-only training network (orange) and a random noise pretrained network (blue). The diagonal line indicates perfect calibration, where confidence matches expected accuracy. (e-f) Averaged confidence and accuracy measured across training epochs for the training dataset. (e) Random noise pretrained network. (f) Data-only training network.
@@
To further validate the effect of pre-calibration, we examined the learning dynamics of networks during subsequent training. To assess the impact of initial loss reduction — where the loss levels of networks were already lowered through random noise pretraining before learning from data — we investigated the training trajectory in a two-dimensional plane of accuracy and loss (Fig. 5a). Notably, the trajectory differed significantly between networks with and without random noise pretraining. Specifically, when comparing the points on both curves with the same accuracy value, we found that the randomly trained network had significantly lower loss values than the network trained solely on data. By selecting arbitrary points of the same accuracy (Fig. 5a, (1) and (2)), we examined the reliability diagram and calibration error for the two networks. We found that networks with random pretraining achieve better confidence calibration than those without pretraining, even with the same accuracy (Fig. 5b, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). Similarly, by selecting arbitrary points of the same total training epochs (Fig. 5a, (3), (4)), we observed that pretrained networks achieve better confidence calibration (Fig. 5c, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). These results demonstrate that random noise pretraining modulates the dynamics of subsequent learning, particularly by reducing the loss level for the same accuracy.
@@
We next examined the relationship between confidence and accuracy throughout the entire training process, using a two-dimensional scatter plot of accuracy versus confidence (Fig. 5d). We found that in pretrained networks, confidence and accuracy were well-aligned, consistently lying near the diagonal line throughout training, while networks without pretraining did not exhibit this alignment (Fig. 5e). Specifically, a significant disparity between confidence and accuracy was initially observed due to early miscalibration, but this gap quickly diminished during random noise pretraining and ultimately disappeared. These initially aligned confidence and accuracy levels facilitate learning with synchronized confidence and accuracy during subsequent data training (Fig. 5f). Thus, calibrating confidence to the chance level through random noise pretraining before learning from data ensures proper calibration of predicted probabilities.
@@
Fig. 6: Out-of-distribution detection using calibrated network confidence
@@
(a) Conceptual illustration of In-Distribution (ID) and Out-Of-Distribution (OOD) samples. ID samples are those that the model has been trained on, while OOD samples are from a different distribution, not seen during training. (b) CIFAR-10 is used as ID data to train the neural network, and SVHN is used as OOD data, which the network has not been trained on. (c) Histogram of confidence for unseen OOD data (left). The distribution is also displayed as a boxplot (right). The horizontal line represents the chance level, indicating ideally calibrated confidence for unknown data, where the model exhibits maximized uncertainty. (d) Schematic of the OOD detection task. After training the network on ID data, both ID and OOD samples are presented to the network, and confidence is measured to determine whether the samples belong to the ID or OOD category. (e) Measured confidence for ID and OOD data and their Cumulative Distribution Functions (CDFs) in a data-only training network (left) and a random noise pretrained network (right). (f) Performance of the OOD detection task using network confidence. (Left) Receiver Operating Characteristic (ROC) curve, where the diagonal line represents random classification, and the point (0, 1) represents an ideal classifier. (Right) Performance of the OOD detection task, measured by the area under the ROC curve (AUROC).
@@
Lastly, we investigated the effect of random noise pretraining on confidence calibration for out-of-distribution (OOD) data. OOD samples refer to unseen data that the model has not been trained on, in contrast to in-distribution (ID) samples, which the model has been trained on (Fig. 6a). After training our model on CIFAR-10 (ID), we measured the network’s response to unseen SVHN data (OOD) (Fig. 6b). Notably, networks without random noise pretraining exhibited high confidence, even though they had no knowledge of OOD data (Fig. 6c, Left). In contrast, randomly pretrained networks showed significantly lower confidence for OOD samples, aligning closer to the chance level, which represents the ideal confidence for ”unknown” data. We confirmed that randomly pretrained networks exhibited lower average confidence, closer to the chance level, for OOD samples (Fig. 6c, Right; w/ vs. w/o random pretraining, $n_{\text{Sample}}=10000$, Wilcoxon rank-sum test, $P<10^{-3}$).
@@
We further investigated how calibrated confidence, for both in-distribution (ID) and out-of-distribution (OOD) samples, enhances the OOD detection task. To do this, we implemented a simple framework that used only raw confidence values to distinguish between ID and OOD samples (Fig. 6d). We analyzed the confidence histograms of ID and OOD samples and their cumulative distribution functions (Fig. 6e). In networks without pretraining, we observed significant overlap in the confidence histograms between ID and OOD samples, making it challenging to differentiate between the two groups (Fig. 6e, Left). In contrast, networks with random noise pretraining showed less overlap, making the task easier (Fig. 6e, Right). We then measured the classification accuracy at various thresholds and plotted the results in the receiver operating characteristic (ROC) curve (Fig. 6f, Left). By calculating the area under the ROC curve (AUROC), we quantified the performance of the networks (Fig. 6f, Right). The results indicate that random noise pretraining significantly improves OOD detection performance (Fig. 6f, Right; w/ vs. w/o random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). These findings suggest that calibrated confidence helps maintain low confidence for unseen OOD data even after training with ID data. This implies that properly calibrated confidence, for both ID and OOD data, enables robust discrimination between ”known” and ”unknown” data without the need for post-processing or auxiliary computations.
@@ Discussion
We demonstrated that pretraining with random noise enables neural networks to learn calibrated probabilities, aligning their confidence levels with expected accuracy. Specifically, we found that random noise training maximizes the initial uncertainty of the network’s predictions, calibrating its confidence to the chance level. We validated that this pre-calibration allows networks to learn with reduced calibration errors. Furthermore, we showed that the network can effectively detect previously unseen data, i.e., out-of-distribution samples. These results suggest that properly calibrating the initial conditions through random noise training can significantly influence learning dynamics and enhance performance.
A longstanding question in machine learning concerns the proper initialization of neural networks for efficient learning. If the initialization is not appropriate, the network may converge to a local minimum or even diverge 55, 56. Previous studies have shown that altering the initialization can help the network converge with higher accuracy 53, 57. Several optimized initialization techniques have been proposed based on the network’s activation function, which were considered optimal starting points for achieving higher accuracy. However, our results suggest that these previously proposed initialization methods are not ideal for learning in terms of calibration, as they tend to be initially overconfident and biased toward specific output neurons. We have confirmed that this initial miscalibration makes it challenging to align confidence with accuracy during subsequent training.
Previous research has primarily addressed overconfidence issues through post-calibration, where “calibrated confidence” is recalculated separately from the raw network output 34, 35, 36, 37, 12. In addition to post-calibration, recent studies have explored the potential of Bayesian neural networks to learn calibrated probabilities 38, 39, 40. However, these methods require significantly more computational resources in addition to the original networks. Moreover, detecting out-of-distribution (OOD) samples necessitates auxiliary computations, such as training additional networks or calculating extra scores to identify unseen samples 15, 16, 41, 42, 43, 44. Traditionally, the calibration of in-distribution (ID) and OOD samples has been treated as separate problems, often requiring additional, complex computations beyond the core learning process. In contrast, our proposed method offers a unified solution for both ID and OOD data, without the need for extra resources or post-processing.
While artificial intelligence struggles with calibration issues through post-calibration or auxiliary computations, human intelligence naturally exhibits meta-cognition — the ability to distinguish between what is known and what is unknown. Previous studies have demonstrated that humans possess an inherent ability to assess probabilistic confidence 47, and that estimating uncertainty plays a key role in complex human behaviors, particularly in decision-making 45, 46, 48. Our proposed method offers a straightforward solution to confidence calibration, enabling human-like uncertainty estimation and fostering meta-cognition. We anticipate that this approach will help address other challenging issues in machine learning caused by calibration, such as machine hallucination in large language models (LLMs).
@@
Modern deep learning models are becoming increasingly complex to enhance their learning capacity. However, over-parameterization often results in miscalibration problems, as highlighted in previous research 12. We discovered that dataset size also plays a crucial role in calibration quality — when the dataset is insufficient, large and complex models experience significant calibration errors. This raises the question of whether an extremely large dataset is required to achieve optimal calibration as model complexity continues to increase. Our proposed method offers a solution by decoupling miscalibration issues from both dataset size and model complexity. This approach provides a simple yet effective strategy for modern deep neural networks as they evolve with more parameters.
We would like to emphasize that our work on pretraining with random noise is inspired by the prenatal developmental process of biological brains. Unlike artificial neural networks, which rely solely on training with real data, biological neural networks undergo preparatory learning before encountering sensory inputs from the external world. This preparation takes place through prenatal and spontaneous neural activity 58, 59, 60, 61. Our recent work has shown that this random noise pretraining process, when combined with feedback alignment algorithms, offers a straightforward strategy that facilitates fast and robust learning without the need for weight transport 49. In the current study, our findings suggest that pretraining with random noise also enables human-like uncertainty calibration. This insight may offer a new perspective on the functional benefits of prenatal development in biological neural networks.
@@
In summary, we conclude that pretraining with random noise effectively pre-calibrates network uncertainty, enabling subsequent learning with aligned accuracy and confidence. These results highlight the importance of the initial state of neural networks for effective data learning 55, 53, 62, 63, 64, 65, as well as the significance of the early developmental process before encountering data 66, 49. Our findings provide new insights into uncertainty calibration, offering potential solutions to challenges arising from calibration issues, such as machine hallucination.
@@ Methods
We employed a multi-layer feedforward model as a simplified representative of conventional neural networks to investigate the miscalibration problem and the effect of random noise pretraining. The model consists of multiple non-linear layers, each performing a series of operations: multiplying weights, adding biases, and passing through a ReLU activation function. To mitigate overfitting, batch normalization was applied to every layer. The hidden layer size was fixed at 256 neurons, with the layer depth varied across different experiments. We observed that our results were fairly consistent regardless of hyperparameter settings, such as the hidden layer size, under typical conditions. The network weights were randomly initialized following a Gaussian distribution, with a mean of 0 and a standard deviation determined to control the gain across layers, as per the standard initialization method 53. The biases were initialized to zero.
@@
Subsequental training with real data
After pretraining the neural network with random noise, we proceeded to train it with real data. We used a subset of the CIFAR-10 dataset, which consists of 32×32×3 natural images representing 10 categories: airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks. In the primary analysis, we used a training subset of 1000 samples to assess the robustness of calibration following random noise pretraining under challenging conditions. To investigate the effects of training data size and model capacity, we employed feedforward networks with depths of 2, 3, 4, 5, and 6 and trained them on CIFAR-10 subsets of sizes 500, 1000, 2000, 4000, 8000, 16000, and 32000.
@@
Additionally, we analyzed the neural network training using a feedback alignment algorithm, as a biologically plausible alternative to backpropagation that does not require weight transport. In feedback alignment, random synaptic feedback is used to compute the error signal, instead of using forward weights. Results obtained using this method are presented in the supplementary material.
@@
To evaluate the calibration of the model, we used a reliability diagram, which visualizes the relationship between the model’s confidence and its actual accuracy. In this approach, we bin the model’s confidence and compute the expected accuracy within each bin. The reliability diagram then plots the expected accuracy as a function of the model’s confidence.
@@
, where $M$ is the total number of bins, $B_{m}$ is the number of samples in the $m$-th bin, $N$ is the total number of samples, $\text{acc}(B_{m})$ is the accuracy of the $m$-th bin, and $\text{conf}(B_{m})$ is the average confidence in the $m$-th bin. In our simulations, we used 10 bins. To quantitatively measure the calibration of the model, we used the Expected Calibration Error (ECE), which calculates the average difference between the model’s confidence and its actual accuracy within the binned confidence. ECE is calculated as:
@@
As a proof-of-concept, we first employed a simple toy-model network with a two-dimensional (2D) input space. The input space is defined as the span of two features, each varying between -1 and 1. This allows us to visualize the input space in the 2D plane, where the x-axis represents the first feature and the y-axis represents the second feature. The neural network model consists of two layers, with a hidden layer of size 10. It outputs a two-dimensional vector for binary classification. The model is initialized using He initialization.
@@
To extend this analysis, we performed similar experiments with high-dimensional feedforward neural networks. We sampled high-dimensional random noise and measured the softmax probability output of the network. We visualized the model’s softmax outputs for both untrained networks and those pretrained with random noise. During the random noise pretraining of the high-dimensional networks, we also measured the confidence and accuracy for real data, such as CIFAR-10 and SVHN. Additionally, we calculated the expected calibration error (ECE) for CIFAR-10 and SVHN during random noise pretraining.
@@
The datasets used in this study are publicly available: https://www.cs.toronto.edu/~kriz/cifar.html (CIFAR-10 and CIFAR-100), http://ufldl.stanford.edu/housenumbers (SVHN).
@@
The analysis and simulations were performed using Python 3.11 (Python software foundation) with PyTorch 2.1 and NumPy 1.26.0. Statistical tests were conducted using SciPy 1.11.4 was used to perform the statistical test and analysis. The code used in this study is available from the corresponding author upon request.

### v2

@@ Results
We further investigated whether the pre-calibration effect observed with random noise training extends to unseen, more realistic data. Specifically, we examined how the network’s confidence on the CIFAR-10 and SVHN 45 datasets changes during random noise pretraining (Supplementary Fig. 6a, Left) by measuring the network’s predictions for real data after each epoch of pretraining. Notably, we observed that the confidence for unseen CIFAR-10 data was properly calibrated throughout the random noise training phase, approaching the chance level, while accuracy remained at chance level (Supplementary Fig. 6a, Right). We confirmed that this trend was consistent across different unseen datasets, including CIFAR-10 (Supplementary Fig. 6b, CIFAR-10; w/ vs. w/o random pretraining, $n_{\text{trial}}=10000$, Wilcoxon rank-sum test, $P<10^{-3}$) and SVHN (Supplementary Fig. 6b, SVHN; w/ vs. w/o random pretraining, $n_{\text{trial}}=10000$, rank-sum test, $P<10^{-3}$). The network maximized uncertainty for unseen data, with accuracy remaining at the chance level, thereby aligning confidence with accuracy. As a result, calibration error — the discrepancy between confidence and accuracy — was significantly reduced for unseen data (Supplementary Fig. 6c, w/ vs. w/o random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, CIFAR-10, $P<10^{-3}$; SVHN, $P<10^{-3}$). These findings demonstrate that conventionally initialized networks tend to exhibit overconfidence and class bias even before training on data, but random noise pretraining effectively eliminates this initial bias and calibrates confidence to the chance level, maximizing uncertainty for unseen datasets. This suggests that random noise pretraining can serve as a novel initialization strategy, offering an alternative to conventional random initialization(He et al., 2015).
Aligning confidence levels with actual accuracy
Fig. 4: Pre-calibration enables learning with matching confidence and accuracy.
@@
(a) Visualization of loss-accuracy duality. Each dot represents the test loss and accuracy at each epoch of training. Four points are selected for calibration comparison: networks achieving the same accuracy for both the random noise-pretrained network and the data-only training network (①, ②), as well as fully converged networks trained for the same number of epochs (③, ④). (b-c) Comparison between random noise-pretrained networks and data-only training networks. (Left) Reliability diagram, showing the relationship between predicted confidence and accuracy. (Right) Calibration error, measured as Expected Calibration Error (ECE) (b) Accuracy-controlled comparison (① vs. ②). (c) Epoch-controlled comparison (③ vs. ④). (d) Scatter plot of confidence versus accuracy for the training dataset in a data-only training network (orange) and a random noise pretrained network (blue). The diagonal line indicates perfect calibration, where confidence matches expected accuracy. (e-f) Averaged confidence and accuracy measured across training epochs for the training dataset. (e) Data-only training network. (f) Random noise-pretrained network.
@@
To better understand the impact of pre-calibration, we examined the learning dynamics of networks during data training, by analyzing the training trajectory in a two-dimensional plane of accuracy and loss (Fig. 4a). The trajectory showed significant differences between networks with and without random noise pretraining, primarily due to the initial loss reduction during pretraining. Specifically, for points on both curves with the same accuracy, the randomly trained network exhibited significantly lower loss than the network trained solely on data. By selecting arbitrary points of identical accuracy (Fig. 4a, ① and ②), we analyzed the reliability diagram and calibration error for both networks. We found that networks with random pretraining had better confidence calibration compared to those without pretraining, even with the same accuracy (Fig. 4b, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). Similarly, when selecting arbitrary points with the same total training epochs (Fig. 4a, ③, ④), pretrained networks again demonstrated better confidence calibration (Fig. 4c, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). These results highlight how random noise pretraining influences the dynamics of subsequent learning, particularly by reducing initial loss.
@@
We next explored the relationship between confidence and accuracy throughout the entire training process (Fig. 4d). In pretrained networks, confidence and accuracy remained well-aligned, consistently following the diagonal line during training (Fig. 4d, w/). In contrast, networks without pretraining did not exhibit this alignment (Fig. 4d, w/o and Fig. 4e). Notably, a significant disparity between confidence and accuracy was initially observed due to early miscalibration, but this gap quickly diminished and disappeared during random noise pretraining (Fig. 4d, Random). This alignment of confidence and accuracy supports more synchronized learning during subsequent data training (Fig. 4f). Thus, calibrating confidence to the chance level through random noise pretraining ensures proper confidence calibration.
@@
Fig. 5: Out-of-distribution detection using calibrated network confidence.
@@
(a) Conceptual illustration of In-Distribution (ID) and Out-Of-Distribution (OOD) samples. ID samples are those that the model has been trained on, while OOD samples are from a different distribution, not seen during training. CIFAR-10 is used as ID data to train the neural network, and SVHN is used as OOD data, which the network has not been trained on. (b) Histogram of confidence for unseen OOD data (left). The distribution is also displayed as a boxplot (right). The horizontal line represents the chance level, indicating ideally calibrated confidence for unknown data, where the model exhibits maximized uncertainty. (c) Schematic of the OOD detection task. After training the network on ID data, both ID and OOD samples are presented to the network, and confidence is measured to determine whether the samples belong to the ID or OOD category. (d) Performance of the OOD detection task using network confidence. (Left) Receiver Operating Characteristic (ROC) curve, where the diagonal line represents random classification, and the point (0, 1) represents an ideal classifier. (Right) Performance of the OOD detection task, measured by the area under the ROC curve (AUROC).
@@
Finally, we examined the effect of random noise pretraining on detecting out-of-distribution (OOD) data. OOD samples refer to “unseen” data the model has not been trained on, while in-distribution (ID) samples are those the model has been trained to recognize (Fig. 5a). After training on CIFAR-10 (ID), we evaluated the model’s response to unseen SVHN data (OOD). Networks without random noise pretraining exhibited high confidence on OOD samples, even though they had no prior knowledge of them (Fig. 5b, Left). In contrast, networks with random pretraining showed significantly lower confidence on OOD samples, aligning closer to the chance level — ideal for “unknown” data. We confirmed that randomly pretrained networks had lower average confidence for OOD samples, closer to chance level (Fig. 5b, Right; w/ vs. w/o random pretraining, $n_{\text{sample}}=10000$, Wilcoxon rank-sum test, $P<10^{-3}$).
@@
We further explored how calibrated confidence for both in-distribution (ID) and out-of-distribution (OOD) samples can improve OOD detection. To achieve this, we implemented a simple framework where raw confidence values were used to distinguish ID from OOD samples (Fig. 5c), with samples below an arbitrary threshold being classified as OOD. We then assessed the OOD detection performance across various thresholds and used these results to plot the receiver operating characteristic (ROC) curve (Fig. 5d, Left). To quantify performance, we calculated the area under the ROC curve (AUROC) (Fig. 5d, Right) and found that random noise pretraining significantly improved OOD detection (Fig. 5d, Right; w/ vs. w/o random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$). These results suggest that calibrated confidence maintains low confidence for unseen OOD data, enabling robust discrimination between “known” and “unknown” data without the need for post-processing or additional computations.
@@ Discussion
The ability to estimate uncertainty and recognize the unknown is crucial for real-world intelligent systems. Human intelligence naturally exhibits meta-cognition — the ability to assess probabilistic confidence 29, 33 and distinguish between what is known and unknown. This capacity plays a crucial role in complex human behaviors, particularly in decision-making 30, 31, 32. Without this meta-cognitive ability to estimate uncertainty, intelligent systems may make critical errors in real-world scenarios where incorrect decisions can be costly 14, 15. Despite recent advances in performance 2, 3, 4, 5, deep neural networks continue to struggle with uncertainty calibration 9, 10, 11, as our findings have demonstrated. This issue poses a significant challenge for integrating artificial intelligence into daily life, despite its advanced capabilities.
Hallucinations in large language models (LLMs) 16, 17 are a prominent example of how miscalibrated uncertainty can make artificial intelligence systems unreliable. While LLMs exhibit expert-level proficiency across various domains, such as science, programming, and healthcare, they often produce incorrect answers with high confidence, making it difficult to trust their outputs in critical situations 18, 19. Beyond LLMs, miscalibration has long been a challenge in applying AI across real-world domains. For instance, although autonomous vehicle prototypes emerged over a decade ago, their integration into daily life remains difficult due to reliability concerns, highlighted by a fatal accident caused by the control system misinterpreting an object 46. Similar issues arise in healthcare, where models that classify diseases from radiological images often make overconfident but incorrect predictions when presented with outlier inputs 47. Despite achieving expert-level accuracy on medical board exams, LLMs frequently fail to recognize their knowledge limitations and provide incorrect yet confident answers 48. Relying on such overconfident judgments in medical procedures could jeopardize patient safety. These examples highlight how miscalibrated predictions can lead to catastrophic decision-making errors, emphasizing the need to address miscalibration issues before AI is deployed in critical applications.
@@
Our results demonstrate that these miscalibrations of uncertainty are inevitable in modern deep neural networks as they become more complex to enhance their learning capacity 2, 3, 4, 5. We found that both dataset size and model complexity systematically influence calibration quality — when the dataset is insufficient, large and complex models experience significant calibration errors. This raises the question of whether an extremely large dataset is always necessary for achieving optimal calibration as model complexity increases. However, acquiring high-quality training data is often costly in the real world, and in some domains, the available data is inherently limited, such as text data in minority languages or medical data on rare diseases. Given these practical challenges, relying solely on expanding dataset size is not a feasible solution as model complexity grows. Our proposed method addresses this by decoupling miscalibration issues from both dataset size and model complexity, offering a simple yet effective strategy for modern deep neural networks as they evolve with increasing parameters.
A longstanding question in machine learning is how to properly initialize neural networks for efficient learning, as the network may converge to a local minimum or diverge depending on the initial condition 49, 50. Previous studies have shown that adjusting initialization can help networks converge with higher accuracy 44, 51, leading to the development of optimized techniques for initialization, a common starting point for training. However, our findings suggest that these traditional methods may not be ideal for learning due to issues with confidence calibration — networks initialized by conventional methods tend to suffer from overconfidence and biases toward specific outputs. The widely adopted practice of random initialization in deep learning is actually a key factor contributing to the miscalibration of network uncertainty; however, pretraining with random noise offers a straightforward and effective solution, thereby proposing a new approach for network initialization. This approach is applicable to all types of network models, and remains effective regardless of model size, complexity, or other characteristics. Unlike previous uncertainty calibration methods that rely on auxiliary computation or post-processing 9, 26, 27, 28, 22, 23, 24, 25, our findings demonstrate that simply altering the initialization of learning through random noise pretraining significantly changes learning dynamics, aligning confidence with accuracy.
Our work on pretraining with random noise is inspired by the prenatal developmental process of biological brains. While spontaneous neural activity in the brain 40, 41, 38, 39 is recognized as important for development, its functional role remains unclear. Recently, artificial neural networks have been considered as a powerful framework for understanding biological neural processing 52, 53, 54. Our results not only provide an effective and practical solution to calibration issues in machine learning models, but also offer insights into biological prenatal processes. In addition to employing the backpropagation algorithm 42, we explored the benefits of random noise pretraining within biologically plausible learning rules, such as feedback alignment 43. Interestingly, we observed severe miscalibration when using local learning rules, but random noise pretraining improved calibration substantially, consistent with our primary findings. Our recent research also indicated that initial calibration and conditioning through random noise pretraining can significantly enhance the learning efficiency of biologically plausible learning systems 34. These findings suggest that the pre-sensory learning process in the biological brain may involve learning the “unknown” before acquiring knowledge through sensory experience, potentially marking a key developmental step toward meta-cognition.
While the current study provides valuable insights, it is important to acknowledge its limitations. Our model was not directly validated through real-world applications or benchmarking, and the findings are based on relatively simple tasks and smaller-scale networks compared to state-of-the-art models, such as large language models (LLMs). Given that training a single LLM can incur millions of dollars in computational costs, it is often impractical to systematically investigate their calibration in a typical laboratory setting. However, we are confident that the results obtained in this study can be replicated in more complex, real-world models, whether in image classification, text generation, or other domains. A key characteristic across various architectures and task domains is that the final outputs are probability distributions. Our pretraining method remains broadly applicable and effective as long as the model generates probability distributions for final decision-making, regardless of its structure, scale, or domain. Moving forward, applying our model to a range of commercial AI systems to further validate and refine these results presents a valuable avenue for future research.
@@
In summary, we conclude that pretraining with random noise effectively pre-calibrates network uncertainty, enabling subsequent learning with aligned accuracy and confidence. These results highlight the importance of the initial state of neural networks for effective data learning 55, 56, 57, 58, 49, 44, as well as the significance of the early developmental process before encountering data 59, 34. Our findings provide new insights into uncertainty calibration, offering potential solutions to challenges arising from calibration issues, such as machine hallucination.
@@ Methods
We employed a multi-layer feedforward model as a simplified representative of conventional neural networks to investigate the miscalibration problem and the effect of random noise pretraining. The model consists of multiple non-linear layers, each performing a series of operations: multiplying weights, adding biases, and passing through a ReLU activation function. To mitigate overfitting, batch normalization was applied to every layer. The hidden layer size was fixed at 256 neurons, with the layer depth varied across different experiments. The network weights were randomly initialized following a Gaussian distribution, with a mean of 0 and a standard deviation determined to control the gain across layers, as per the standard initialization method 44. The biases were initialized to zero. We observed that our results were fairly consistent regardless of hyperparameter settings, such as the hidden layer size, under typical conditions.
@@
Subsequent training with real data
After pretraining the neural network with random noise, we proceeded to train it with real data. We used a subset of the CIFAR-10 35 dataset, which consists of 32×32×3 natural images representing 10 categories: airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks. In the primary analysis, we used a training subset of 1000 samples to assess the robustness of calibration following random noise pretraining under challenging conditions. To investigate the effects of training data size and model capacity, we employed feedforward networks with depths of 2, 3, 4, 5, and 6 and trained them on CIFAR-10 subsets of sizes 500, 1000, 2000, 4000, 8000, 16000, and 32000.
@@
Additionally, we analyzed the neural network training using a feedback alignment 43 algorithm, as a biologically plausible alternative to backpropagation that does not require weight transport. In feedback alignment, random synaptic feedback is used to compute the error signal, instead of using forward weights. Results obtained using this method are presented in the supplementary material.
@@
To evaluate the calibration of the model, we used a reliability diagram 36, 22, which visualizes the relationship between the model’s confidence and its actual accuracy. In this approach, we bin the model’s confidence and compute the expected accuracy within each bin. The reliability diagram then plots the expected accuracy as a function of the model’s confidence
@@
where $M$ is the total number of bins, $B_{m}$ is the number of samples in the $m$-th bin, $N$ is the total number of samples, $\text{acc}(B_{m})$ is the accuracy of the $m$-th bin, and $\text{conf}(B_{m})$ is the average confidence in the $m$-th bin. In our simulations, we used 10 bins.
To quantitatively measure the calibration of the model, we used the Expected Calibration Error (ECE) 37, which calculates the average difference between the model’s confidence and its actual accuracy within the binned confidence. ECE is calculated as:
@@
As a proof-of-concept, we first employed a simple toy-model network with a two-dimensional (2D) input space. The input space is defined as the span of two features, each varying between -1 and 1. This allows us to visualize the input space in the 2D plane, where the x-axis represents the first feature, and the y-axis represents the second feature. The neural network model consists of two layers, with a hidden layer of size 10. It outputs a two-dimensional vector for binary classification. The model is initialized using He initialization.
@@
To extend this analysis, we performed similar experiments with high-dimensional feedforward neural networks. We sampled high-dimensional random noise and measured the SoftMax probability output of the network. We visualized the model’s SoftMax outputs for both untrained networks and those pretrained with random noise. During the random noise pretraining of the high-dimensional networks, we also measured the confidence and accuracy for real data, such as CIFAR-10 and SVHN 45. Additionally, we calculated the Expected Calibration Error (ECE) for CIFAR-10 and SVHN during random noise pretraining.
@@
The datasets used in this study are publicly available: https://www.cs.toronto.edu/~kriz/cifar.html (CIFAR-10), http://ufldl.stanford.edu/housenumbers (SVHN).
@@
The analysis and simulations were performed using Python 3.11 (Python software foundation) with PyTorch 2.1 and NumPy 1.26.0. Statistical tests were conducted using SciPy 1.11.4 was used to perform the statistical test and analysis. The code used in this work is available at https://github.com/cogilab/Random2.